mathslover (mathslover):

Given : \(x^2-x-1=0\) then : \(x^8 + \frac{1}{x^8}=?\)

OpenStudy (anonymous):

have u tried it yet?

mathslover (mathslover):

For better visual : \[\large{x^2-x-1=0}\] \[\large{x^8+\frac{1}{x^8}=?}\] Yes mukushla.... I tried 1) put x = \(\frac{1}{x-1}\) but that is too long...

mathslover (mathslover):

@mukushla @hartnn @sauravshakya @ParthKohli

OpenStudy (anonymous):

Cant we find the value of x and substitute it

mathslover (mathslover):

mentioned above but too long...

mathslover (mathslover):

Well let me know who can solve this I know the correct soln .... but I will like to know who knows it

mathslover (mathslover):

Type the answer here

OpenStudy (anonymous):

\[x^2-1=x\]\[x-\frac{1}{x}=1\]square both sides 4 time

mathslover (mathslover):

@allstudentshere : put your answer here

Parth (parthkohli):

\[x^2 = x +1\]\[x = \sqrt{x + 1}\] \[x^8 = (x + 1)^4\]\[{1 \over x^8} = {1 \over (x + 1)^4}\]Hmm...

mathslover (mathslover):

:) only answer friends

mathslover (mathslover):

1st guess made by mukushla is "wrong" sorry

OpenStudy (anonymous):

lol 47

OpenStudy (anonymous):

Yep 47

OpenStudy (anonymous):

But that method was not too long

OpenStudy (anonymous):

Solve x^2-x-1=0 u will get 

OpenStudy (anonymous):

And substitute it in x^8+1/x^8

mathslover (mathslover):

\[\large{x^8 + \frac{1}{x^8}}\] \[\large{x^8+\frac{1}{x^8}+2-2}\] \[\large{(x^4+\frac{1}{x^4})^2-2}\] \[\large{[(x^2+\frac{1}{x^2})^2-2]^2}\] like this

mathslover (mathslover):

Guys: mukushla and saurav have right answers

OpenStudy (anonymous):

Oh........ But @mukushla method is much better

hartnn (hartnn):

i did the same thing, still i got -1....maybe i took, -2 instead of +2 somewhere...

mathslover (mathslover):

\[\large{[[(x+\frac{1}{x})^2-2]^2-2]^2-2}\] \[\large{\frac{1}{x}=-(\frac{\sqrt{5}+1}{2})}\]

mathslover (mathslover):

did this...

Parth (parthkohli):

Again I ask, where did you get +2 and -2 from?

mathslover (mathslover):

\[\large{x=\frac{-2}{\sqrt{5}+1}}\] \[\large{\textbf{Put this value in the formula we willl get 47 as answer}}\]

Parth (parthkohli):

No, I don't mean that. Why did you add and subtract 2 and -2 respectively?

mathslover (mathslover):

Oh k : \[\large{(x+\frac{1}{x})^2 = (x^2+\frac{1}{x^2}+2)}\] hence to make that as look like this... we did subtract 2 and -2 ..

Parth (parthkohli):

Which formula did you use to complete the square?

OpenStudy (anonymous):

x^2-x-1=0 x(x-1-1/x)=0 x-1/x-1=0 x-1/x=1 x^2-2*x*1/x +1/x^2 =1 x^2+1/x^2=3 x^4+1/x^4+2=9 x^4+1/x^4=7 x^8+1/x^8+2=49 x^8+1/x^8=47 @mukushla mentioned this method

OpenStudy (anonymous):

\[x^2-x-1=0\] \[x^{16}-kx^8+1=0\] \[x^8=y\] \[y^2-ky+1=0\] \[y=\frac{k \pm \sqrt{k^2-4}}{2}=x^8\]

OpenStudy (anonymous):

\[\frac{k \pm \sqrt{k^2-4}}{2}=(\frac{1 \pm \sqrt{1+4}}{2})^8\]

mathslover (mathslover):

1+4 = 5

OpenStudy (anonymous):

@mathslover the answer is 47 right?

mathslover (mathslover):

Correct saurav sir :)

OpenStudy (anonymous):

And u got it?

mathslover (mathslover):

Yep.... I knew the soln (and answer) but just wanted to know another way

OpenStudy (lgbasallote):

ahh it's so beautiful to see so many lovers of math/s

mathslover (mathslover):

including you :) --> lgba --> maths lover

OpenStudy (lgbasallote):

..............

OpenStudy (lgbasallote):

don't ever say that. it disgusts me.

mathslover (mathslover):

I will like to say again though I care about your feelings.... sorry lgba :'()